OpenStudy (anonymous):

Which of the following equations represents a perpendicular line to 6x - 2y = 20 y = 3x + 2 y = -3x + 4 y = 1/3x-3 y= -1/3x-5

OpenStudy (anonymous):

convert 6x-2y=20 to y=mx+b form :)

OpenStudy (anonymous):

so y = 6x-2y?

OpenStudy (anonymous):

y=3x+10 :)

OpenStudy (anonymous):

so what now

OpenStudy (anonymous):

now the opposite reciprocal of m is the slope of the new line...so the slope is -1/3 so any equation with a slope of -1/3 is perpendicular to 6x-2y=20 :)

OpenStudy (anonymous):

so the last one :)

OpenStudy (anonymous):

yup yup! ;)

OpenStudy (anonymous):

k hope i helped! byee! :)